# On the influence of electrolytic gradient orientation on phoretic transport in dead-end pores

### v1

@@ III.A.1 Persistence time
$\frac{\bm{u}_{p}}{\bm{u}_{p}\left.\right|_{(t=0)}}=\begin{cases}\displaystyle\frac{\frac{4}{\pi}\sin\left(\frac{\pi x}{2}\right)}{\frac{4}{\pi}\sin\left(\frac{\pi x}{2}\right)+\beta e^{\lambda t}}\sim\frac{1}{1+\beta e^{\lambda t}}\;\text{for}\;\beta\ll 1\\[12.0pt] \displaystyle\frac{\frac{4}{\pi}\sin\left(\frac{\pi x}{2}\right)-1}{\frac{4}{\pi}\sin\left(\frac{\pi x}{2}\right)-e^{\lambda t}}\sim\frac{1}{e^{\lambda t}}\;\text{for}\;\beta\gg 1\end{cases}$ (3)
The persistence of extreme solute-in mode ($\beta\gg 1$) shows independence with respect to the magnitude of applied solute gradient, whereas the extreme solute-out mode ($\beta\ll 1$) demonstrates a delayed exponential decay that scales with $\sim\ln\left[\beta^{-1/\lambda}\right]$, where $1/\lambda=4/\pi^{2}\approx 0.4$. The latter is also explicitly shown in Fig. 4 (b), where a reduction in $\beta$ delays the exponential decay of phoretic velocity. We note that, the only condition that can sustain the diffusiophoretic process indefinitely is an environment absolutely free of solute ($\beta=0$) at all times.
III.A.2 Effectiveness
Persistence time characterization aids in concluding that solute-out mode has relatively larger lifespan. As an additional performance metric, we evaluate the ‘effectiveness’, which quantifies the total colloids withdrawn in the presence of solute gradients relative to bare colloidal diffusion (in the absence of solute gradient, i.e., $\beta=1$). For injection, it is defined as the ratio of total colloids injected in the presence of solute gradients relative to bare colloidal diffusion.
@@ III.A.2 Effectiveness
III.B Withdrawal and injection of colloids with variable mobility
We now extend our analysis by relaxing the assumption of constant mobility. Specifically, colloids’ zeta potential is the function of the local solute concentration Kirby and Hasselbrink Jr [2004], influencing the concentration-dependent mobility, i.e., the mobility experiences spatio-temporal variation. The mobility dependence on concentration is reported to arise due to counterion absorption into the Stern layer. The consequent rise in counterion shielding is associated with alteration in the debye layer thickness, and yields in the modification of zeta potential. The ‘constant charge model’ Kirby and Hasselbrink Jr [2004] assumes that the surface charge density remains fixed and for base zeta potentials larger than thermal potential, its modification is logarithmic $\left(\sim a+b\log(c)\right)$, which would consequently alter the mobility $\Gamma_{\text{p}}(c)$ as:
$\frac{\varepsilon}{2\mu}\left(\frac{k_{\text{B}}T}{ze}\right)^{2}\left[2\beta_{s}\frac{Ze\zeta_{p}(c)}{k_{B}T}+8\ln cosh\left(\frac{Ze\zeta_{p}(c)}{4k_{\text{B}}T}\right)\right],$ (5)
where $\beta_{s}(=\frac{\mathcal{D}_{+}-\mathcal{D}_{-}}{\mathcal{D}_{+}+\mathcal{D}_{-}})$, $\varepsilon$, $\mu$, $k_{B}$, $z$, $e$, and $T$ represent the diffusivity contrast between cation and anion, medium permittivity, medium viscosity, Boltzmann constant, ionic valencies, elementary charge, and absolute temperature, respectively. The expression for the zeta potential is motivated by the study by Akdeniz et al. [2023] for PS-PEG particle, they found $\zeta_{p}=a+b\log_{10}(c^{\ast}/c_{ch})$, where $a=-27.43$ mV, $b=22.63$ mV, and $c^{\ast}$ ranges from $1-10$ mM. In our analysis below, we compare the modifications to colloidal density profile that are caused by concentration-dependent mobility.
@@ III.B Withdrawal and injection of colloids with variable mobility
III.C Influence of wall-induced osmotic flow
@@
Figure 9: Spatiotemporal maps with osmotic flow interference. (a) Withdrawal and (b) injection contours of colloids at $0.2\,t_{d}$ and $1\,t_{d}$. Line plots show variation of y-averaged concentration profile $\langle n\rangle{}_{y}$ with osmotic flow ($\Gamma_{p}=f(c)$, $\Gamma_{w}=g(c)$; solid line) and without osmotic flow ($\Gamma_{w}=0$; dot-dashed line). The inset in solute-out plot show the absence of particle banding when phoretic effects are absent. The plots in the bottom row show effectiveness of (a) withdrawal and (b) injection cases for solute out (red) and solute in (blue) modes. Dot-dashed curves correspond to results for $\Gamma_{w}=0$. Inset shows the intensity of lateral flow velocity near the pore mouth, which tends to push the particles towards the centerline. Parameters: $D_{s}=1.61\times 10^{-9}$ m2/s and $D_{p}=4\times 10^{-13}$ m2/s.
@@
Here we consider the cases when walls exhibit significant mobility such that, in the presence of external solute gradient, a diffusio-osmotic slip is generated Anderson [1989] ($\bm{u}_{\text{slip}}=-\frac{\Gamma_{w}(c)}{D_{s}}\nabla\log c$). This slip-driven flow further renders a bulk motion in the pore and has been shown to affect particle dispersion Shin et al. [2016], Kar et al. [2015]. It has also been utilized to trap colloids and vesicles in bioassays Rasmussen et al. [2020], as well as for performing low-cost zeta potentiometry Shin et al. [2017a], Ault et al. [2019]. To begin the analysis, we first note from Eq.(2) that osmotic flow would not contribute significantly to the colloidal dispersion in $\text{Pe}\ll\Gamma_{p}/D_{s}$ regime. Hence in the discussion below, we consider the wall mobilities that yield osmotic flows in $\text{Pe}\sim\Gamma_{p}/D_{s}$ regime. The chosen parameters yield in osmotic and phoretic velocities that broadly match with the experiments of Kar et al. [2015].
@@
Figures. 8 (a,b) illustrate the horizontal component of osmotic flow that occurs for the two solute-gradient modes. To render the shape of bi-directional osmotic flows similar for the two modes (ensuring an equitable comparison), the wall mobilities ($\Gamma_{w}$) are set such that positive values are assigned for the solute-out mode and negative values for the solute-in mode. Other combinations would yield a bi-directional flow of inverted shape. In Fig. 8(a,b), the wall mobility senses the chemical gradients and induces a strong outward wall slip, and as the net flow rate within the pore is zero, a parabolic inflow develops away from the walls Ault et al. [2019]. For solute-out mode, this bidirectional flow decays monotonically along the pore depth, whereas non-monotonically for solute-in mode; a trend similar to diffusiophoretic velocities shown in 6(a,b). Both of these modes generate a modest convective flow ($\sim\Gamma_{p}/D_{s}$) as shown by the temporal decay of the spatially-averaged flow in §3 in the Supplementary material. Figs. 8 (a,b) depict the vertical component of the osmotic flow, which is responsible for the lateral redistribution of colloids. This flow arises due to the dead-end nature of the pore; since the net flow rate is zero, the slip-driven motion of the incompressible fluid ($\partial\bm{u}_{fx}/\partial x\neq 0$) induces a nonzero $\bm{u}_{fy}$ component. For the chosen signs of wall mobilities (or surface potentials), $\bm{u}_{fy}$ directs the colloids away from the centerline. Interestingly, while recirculation occurs near the pore entry in the solute-out mode (as observed here and reported in ref. Shin et al. [2016]), the solute-in mode instead exhibits recirculation at the end of the pore. This is because the osmotic slip is sustained throughout the pore length and $\partial\bm{u}_{fx}/\partial x$ is largest near $x^{*}=L$, and thus $\bm{u}_{fy}$ is largest here. This difference in recirculation location for the two modes will be shown to cause significant deviations in transport efficacy. Finally, we note that these trends in slip and bulk velocity are exactly reversed between the solute-out and solute-in modes when the mobilities are switched. For completeness, the impact of both combinations of zeta potentials is considered in Fig. 9, which demonstrates the interference of osmotic flow on spatiotemporal patterns of colloids. Note that both combinations yield near-identical effectiveness results.
@@
Fig. 9 (a) shows the results for colloidal withdrawal for two modes (in the top and bottom rows). Here, the osmotic flow is configured such that an outward wall slip ejects particles into the ambient fluid, which is accompanied by an inward bulk flow. The contours are supplemented by the plots that quantify the impact on $y$-averaged profiles. The solute-out mode, which is more responsive to chemical gradients, is substantially influenced by osmotic flow interference, also reflected in the effectiveness measurement shown in Fig. 9(e). For injection, Fig. 9 (b) illustrates similar patterns in concentration profiles, with certain intriguing features that influence both the y-averaged profile and overall effectiveness, a closer look at which follows. The outward osmotic slip at the walls competes with the inward phoretic injection, leading to local colloidal accumulation near the pore entry corners. Whereas, the inward osmotic flow in the channel bulk convects the injecting particles deeper, consequently shaping the injecting band of colloids as a parabola Shin et al. [2016]. In agreement with the observations of Shin et al. [2017a] experiments, we find that this banding is further exacerbated by $\bm{u}_{fy}$ convection, which is strongest near the pore for solute-out and, as shown in Fig. 8(c), acts to push particles toward the outward flow near the walls. As a result of this lateral migration, the effectiveness of the solute-out mode suffers substantially from the osmotic interference—a trend contrary to that of withdrawal.
@@
The results for other combinations of wall mobilities [Fig. 9(c,d)] exhibit similar quantitative trends, with $y$-averaged colloidal concentrations being near-identical to previous case. The spatiotemporal profile, however, now resembles a reverse concentration boundary layer at the top and bottom walls due to the bidirectional flow. Interestingly, the injection profile in Fig. 9 (d) shows inverse banding. This is because, for these wall mobilities, the solute-out mode induces inwawrd osmotic slip and outward bulk flow; the latter competes with the inward phoretic injection, leading to local colloidal accumulation near the central axis of the pore entry. This banding is further shaped by $\bm{u}_{fy}$ convection [shown in Fig. 9(f) inset]. This lateral convection is strongest near the pore for solute-out and, as shown in the inset contour, acts to push particles toward the outward flow near the centerline, which consequently affects the injection efficiency as shown in Fig. 9(f).
IV Conclusions
@@
The current results can aid in a deeper understanding of particle and drug transport in porous biofilms Somasundar et al. [2023] and provide insights into the injection and withdrawal of non-motile microorganisms, recently shown to be influenced by surfactant gradients Doan et al. [2020]. Additionally, while the often-overlooked solute-in mode exhibits slower colloidal injection and withdrawal, its distinct characteristics may be advantageous in applications requiring uniform dispersion and deep penetration Velegol et al. [2016]. For example, this mechanism could be beneficial in shear-thickening fabric manufacturing, where rapid and homogeneous particle infusion into fabric pores can enhance mechanical properties such as tensile strength and durability He et al. [2018], Talreja et al. [2017].

### v2

@@ 0.2-1 $\mu$m polystyrene and
$\frac{\bm{u}_{p}}{\bm{u}_{p}\left.\right|_{(t{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\to}0)}}=\begin{cases}\displaystyle\frac{\frac{4}{\pi}\sin\left(\frac{\pi x}{2}\right)}{\frac{4}{\pi}\sin\left(\frac{\pi x}{2}\right)+\beta e^{\lambda t}}\sim\frac{1}{1+\beta e^{\lambda t}}\;\text{for}\;\beta\ll 1\\[12.0pt] \displaystyle\frac{\frac{4}{\pi}\sin\left(\frac{\pi x}{2}\right)-1}{\frac{4}{\pi}\sin\left(\frac{\pi x}{2}\right)-e^{\lambda t}}\sim\frac{1}{e^{\lambda t}}\;\text{for}\;\beta\gg 1\end{cases}$ (5)
The persistence of extreme solute-in mode ($\beta\gg 1$) shows independence with respect to the magnitude of applied solute gradient, whereas the extreme solute-out mode ($\beta\ll 1$) demonstrates a delayed exponential decay that scales with $\sim\ln\left[\beta^{-1/\lambda}\right]$, where $1/\lambda=4/\pi^{2}\approx 0.4$. The latter is also explicitly shown in Fig. 4 (b), where a reduction in $\beta$ delays the exponential decay of phoretic velocity. We note that, the only condition that can sustain the diffusiophoretic process indefinitely is an environment absolutely free of solute ($\beta=0$) at all times. However, we note that this ideal case cannot be experimentally realized as, at extremely low solute concentrations, the continuum approximation breaks down, as the distance between the ions can become comparable or larger than particle size.
The persistence time can also be analytically estimated within the one-term approximation. Plugging the aforementioned $\bm{u}_{p}$ into $\frac{\bm{u}_{p}}{\bm{u}_{p}|_{t=0.1}}=0.01$, approximately yields the following equation for local persistence time:
$\tau=0.405\log\left[128+126\left(\frac{1-\beta}{\beta}\right)\sin(\pi x/2)\right].$ (6)
This expression can be integrated over the non-dimensional pore length to yield the average persistence $\langle\tau\rangle$, which agrees well with numerical calculation and asymptotic estimates in Fig.4 (a) as empty circles.
Effectiveness
Persistence time characterization aids in concluding that solute-out mode has relatively larger lifespan. As an additional performance metric, we evaluate effectiveness, which, for the case of withdrawal, is the ratio of the total colloids withdrawn in the presence of solute gradients relative to those removed by bare colloidal diffusion (in the absence of solute gradient, i.e., $\beta=1$). For injection, it is defined analogously as the ratio of the total colloids injected in the presence of solute gradients to those injected by pure colloidal diffusion ($\beta=1$).
@@
The mobility model considered here is appropriate for micron-sized colloids, where the Debye layer is extremely thin compared to the particle size. For submicron or nanoscale colloids, however, the Debye layer has a finite thickness that varies with concentration, rendering the mobility a function of the local solute concentration, which evolves both spatially and temporally 43, 40. In the Appendix, we incorporate this mobility model and show that while finite Debye layer effects introduce quantitative modifications—such as dampening of colloidal withdrawal and injection—they do not affect the qualitative conclusions presented here.
@@
Withdrawal and injection of colloids with variable mobility
We now extend our analysis by relaxing the assumption of constant mobility. Specifically, colloids’ zeta potential is the function of the local solute concentration, influencing the concentration-dependent mobility, i.e., the mobility experiences spatio-temporal variation. This variability in mobility arises from counterion shielding of surface charges (associated with changes in the Debye layer thickness), resulting in a concentration-dependent modification of the zeta potential 41. This ‘constant charge model’ assumes that the surface charge density remains fixed and for base zeta potentials larger than thermal potential, its modification is logarithmic $\left(\sim a+b\log(c)\right)$ 41 eq.8, which would consequently alter the mobility $\Gamma_{\text{p}}(c)$ as:
$\frac{\varepsilon}{2\mu}\left(\frac{k_{\text{B}}T}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Z}e}\right)^{2}\left[2\beta_{s}\frac{Ze\zeta_{p}(c)}{k_{B}T}+8\ln cosh\left(\frac{Ze\zeta_{p}(c)}{4k_{\text{B}}T}\right)\right],$ (9)
where $\beta_{s}(=\frac{\mathcal{D}_{+}-\mathcal{D}_{-}}{\mathcal{D}_{+}+\mathcal{D}_{-}})$, $\varepsilon$, $\mu$, $k_{B}$, ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Z}$, $e$, and $T$ represent the diffusivity contrast between cation and anion, medium permittivity, medium viscosity, Boltzmann constant, ionic valencies, elementary charge, and absolute temperature, respectively. The expression for the zeta potential is motivated by the study of Akdeniz et al. 8 on size-controlled PS-PEG particles. They found that $\zeta_{p}=a+b\log_{10}(c^{\ast}/1\text{mM})$, where $a=-27.43$ mV, $b=22.63$ mV, and $c^{\ast}$ ranges from 1 to 10 mM. This expression is valid in the limit where the thickness of the electrical double layer is negligible compared to the particle size—an approximation that holds for the micron-sized particles used in Akdeniz et al. 8. In our analysis below, we compare the modifications to colloidal density profile that are caused by concentration-dependent mobility.
@@
Influence of wall-induced osmotic flow
@@
Figure 9: Spatiotemporal maps with osmotic flow interference. (a) Withdrawal and (b) injection contours of colloids at $0.2\,t_{d}$ and $1\,t_{d}$. Line plots show variation of y-averaged concentration profile $\langle n\rangle{}_{y}$ with osmotic flow ($\Gamma_{p}=f(c)$, $\Gamma_{w}=g(c)$; solid line) and without osmotic flow ($\Gamma_{w}=0$; dot-dashed line). The inset in solute-out plot show the absence of particle banding when phoretic effects are absent. (c,d) Spatiotemporal patterns for other combinations of wall and particle mobilities for withdrawal and injection, respectively. (e,f) Effectiveness of withdrawal (a) and injection (b) cases for solute out (red) and solute in (blue) modes, respectively. Dot-dashed curves correspond to results for $\Gamma_{w}=0$. Inset shows the intensity of lateral flow velocity near the pore mouth, which tends to push the particles towards the centerline. Parameters: $D_{s}=1.61\times 10^{-9}$ m2/s and $D_{p}=4\times 10^{-13}$ m2/s.
@@
Here we consider the cases when walls exhibit significant mobility such that, in the presence of external solute gradient, a diffusio-osmotic slip is generated 15 ($\bm{u}_{\text{slip}}=-\frac{\Gamma_{w}(c)}{D_{s}}\nabla\log c$). This slip-driven flow further renders a bulk motion in the pore and has been shown to affect particle dispersion 2, 1. It has also been utilized to trap colloids and vesicles in bioassays 23, as well as for performing low-cost zeta potentiometry 4, 11. To begin the analysis, we first note from Eq.(2) that osmotic flow would not contribute significantly to the colloidal dispersion in $\text{Pe}\ll\Gamma_{p}/D_{s}$ regime. Hence in the discussion below, we consider the wall mobilities that yield osmotic flows in $\text{Pe}\sim\Gamma_{p}/D_{s}$ regime. The chosen parameters yield in osmotic and phoretic velocities that broadly match with the experiments of Kar et al. 1.
@@
Figures. 8 (a,b) illustrate the horizontal component of osmotic flow that occurs for the two solute-gradient modes. To render the shape of bi-directional osmotic flows similar for the two modes (ensuring an equitable comparison), the wall mobilities ($\Gamma_{w}$) are set such that positive values are assigned for the solute-out mode and negative values for the solute-in mode. Other combinations would yield a bi-directional flow of inverted shape. In Fig. 8(a,b), the wall mobility senses the chemical gradients and induces a strong outward wall slip, and as the net flow rate within the pore is zero, a parabolic inflow develops away from the walls 11. For solute-out mode, this bidirectional flow decays monotonically along the pore depth, whereas non-monotonically for solute-in mode; a trend similar to diffusiophoretic velocities shown in 6(a,b). Both of these modes generate a modest convective flow ($\sim\Gamma_{p}/D_{s}$) as shown by the temporal decay of the spatially-averaged flow in §3 in the Supplementary material. Figs. 8 (a,b) depict the vertical component of the osmotic flow, which is responsible for the lateral redistribution of colloids. This flow arises due to the dead-end nature of the pore; since the net flow rate is zero, the slip-driven motion of the incompressible fluid ($\partial\bm{u}_{fx}/\partial x\neq 0$) induces a nonzero $\bm{u}_{fy}$ component. For the chosen signs of wall mobilities (or surface potentials), $\bm{u}_{fy}$ directs the colloids away from the centerline. Interestingly, while recirculation occurs near the pore entry in the solute-out mode (as observed here and reported in ref. 2, 34), the solute-in mode instead exhibits recirculation at the end of the pore. This is because the osmotic slip is sustained throughout the pore length and $\partial\bm{u}_{fx}/\partial x$ is largest near $x^{*}=L$, and thus $\bm{u}_{fy}$ is largest here. This difference in recirculation location for the two modes will be shown to cause significant deviations in transport efficacy. Finally, we note that these trends in slip and bulk velocity are exactly reversed between the solute-out and solute-in modes when the mobilities are switched. For completeness, the impact of both combinations of zeta potentials is considered in Fig. 9, which demonstrates the interference of osmotic flow on spatiotemporal patterns of colloids. Note that both combinations yield near-identical effectiveness results.
@@
Fig. 9 (a) shows the results for colloidal withdrawal for two modes (in the top and bottom rows). Here, the osmotic flow is configured such that an outward wall slip ejects particles into the ambient fluid, which is accompanied by an inward bulk flow. The contours are supplemented by the plots that quantify the impact on $y$-averaged profiles. The solute-out mode, which is more responsive to chemical gradients, is substantially influenced by osmotic flow interference, also reflected in the effectiveness measurement shown in Fig. 9(e). For injection, Fig. 9 (b) illustrates similar patterns in concentration profiles, with certain intriguing features that influence both the y-averaged profile and overall effectiveness, a closer look at which follows. The outward osmotic slip at the walls competes with the inward phoretic injection, leading to local colloidal accumulation near the pore entry corners. Whereas, the inward osmotic flow in the channel bulk convects the injecting particles deeper, consequently shaping the injecting band of colloids as a parabola 2. In agreement with the observations of Shin et al. 4 experiments, we find that this banding is further exacerbated by $\bm{u}_{fy}$ convection, which is strongest near the pore for solute-out and, as shown in Fig. 8(c), acts to push particles toward the outward flow near the walls. As a result of this lateral migration, the effectiveness of the solute-out mode suffers substantially from the osmotic interference—a trend contrary to that of withdrawal.
@@
The results for other combinations of wall mobilities [Fig. 9(c,d)] exhibit similar quantitative trends, with $y$-averaged colloidal concentrations being near-identical to previous case. The spatiotemporal profile, however, now resembles a reverse concentration boundary layer at the top and bottom walls due to the bidirectional flow. Interestingly, the injection profile in Fig. 9 (d) shows inverse banding. This is because, for these wall mobilities, the solute-out mode induces inwawrd osmotic slip and outward bulk flow; the latter competes with the inward phoretic injection, leading to local colloidal accumulation near the central axis of the pore entry. This banding is further shaped by $\bm{u}_{fy}$ convection [shown in Fig. 9(f) inset]. This lateral convection is strongest near the pore for solute-out and, as shown in the inset contour, acts to push particles toward the outward flow near the centerline, which consequently affects the injection efficiency as shown in Fig. 9(f). Finally, we note that wall-induced diffusioosmosis contributes to the spreading of the colloidal distribution, as evidenced by the $y$-averaged profiles for injections with and without wall slip (solid and dashed curves, respectively) shown in Fig. 9(b,d). Additional results from §4 of the Supplementary Material demonstrate that the extend of spread rendered by the wall mobility of two types ($\Gamma_{w}>0$ and $\Gamma_{w}<0$), consistent with earlier findings 34, 7.
Conclusion
@@
The current results can aid in a deeper understanding of particle and drug transport in porous biofilms 44 and provide insights into the injection and withdrawal of non-motile microorganisms, recently shown to be influenced by surfactant gradients 45. Additionally, while the often-overlooked solute-in mode exhibits slower colloidal injection and withdrawal, its distinct characteristics may be advantageous in applications requiring uniform dispersion and deep penetration 17. For example, this mechanism could be beneficial in shear-thickening fabric manufacturing, where rapid and homogeneous particle infusion into fabric pores can enhance mechanical properties such as tensile strength and durability 46, 47.
Banerjee and Squires 48 devised hydrogel-based ‘soluto-inertial’ source and sink beacons that facilitate spatio-temporal control over colloidal suspensions. Later, for non-electrolytic gradients, Raj et al. 49 modeled 2D colloidal banding and reported optimal designs of such beacons that exhibit time-dependent solute emission/absorption. Recently, Alessio and Gupta 50 modeled pigment cells (chromatophores) as non-electrolytic diffusiophoretic particles and demonstrated that, in the presence of two chromatophore types and two solute species, augmented Turing patterns can emerge with significantly finer length scales. Based on the insights from §III on electrolytic gradients, we hypothesize that colloids with positive mobility will exhibit large and rapid wavefronts when attracted to regions of higher solute concentration, whereas colloids with negative mobility will form slower and uniform patterns. This distinction in pattern morphology and timescales may have important implications for the optimal design of soluto-inertial beacons, colloidal banding phenomena, and the understanding of cellular pattern formation, highlighting the need for further investigation.
Figure A.1: (a) and (d) Particle mobility ($\Gamma_{p}$, m2/s) variation with solute concentration ($c^{\ast}$, mM) inside the pore with and without Debye layer effect for $a=250$ nm and $30$ nm, respectively. (b,c) Particle withdrawal and (e,f) injection patterns for solute-out (solid line) and solute-in mode (dashed line) at $a=250$ nm and $30$ nm, respectively. Parameters: $D_{s}=1\times 10^{-9}$ m2/s, $t=0.5$, $D_{p}=1\times 10^{-12}$ m2/s ($a=250$ nm) and $8.084\times 10^{-12}$ m2/s ($a=30$ nm), $\beta=0.01$ (S-O) and $100$ (S-I).
Figure A.2: (a) and (d) Particle mobility ($\Gamma_{P}$, m2/s) variation along the length with Debye layer effect at $t^{*}=0.2t_{d},0.5t_{d}$ and $1t_{d}$ for $a=250$ nm and $30$ nm, respectively. (b,c) Particle withdrawal and (e,f) injection patterns for solute-out (solid line) and solute-in mode (dashed line) at $a=250$ nm and $30$ nm, respectively. Parameters: $D_{s}=1\times 10^{-9}$ m2/s, $D_{p}=1\times 10^{-12}$ m2/s ($a=250$ nm) and $8.084\times 10^{-12}$ m2/s ($a=30$ nm), $\beta=0.01$ (S-O) and $100$ (S-I).
Figure A.3: (a) and (d) Persistence time variation with applied concentration difference ratio $\beta$ with and without Debye later effect for $a=250$ nm and $30$ nm, respectively. Effectiveness of (b,c) withdrawal and (e,f) injection cases for $a=250$ nm and $30$ nm, respectively. Parameters: $D_{s}=1\times 10^{-9}$ m2/s, $D_{p}=1\times 10^{-12}$ m2/s ($a=250$ nm) and $8.084\times 10^{-12}$ m2/s ($a=30$ nm), $\beta=0.01$ (S-O) and $100$ (S-I).
Appendix
When colloids are of sub-micron & nanometer size, the typical ($\sim 1-50$ nm) Debye layer (DL) can no longer be considered as infinitesimally thin compared to colloid radius. Here we utilize the mobility corrections from 43, 40 to account for the finite DL effects to predict the spatiotemporal patterns and performance. The phoretic mobility is modified as follows:
$\Gamma_{p}=\frac{\varepsilon}{\mu}\left[\frac{k_{b}T}{e}\beta_{s}\zeta\,\Theta_{1}\left(\frac{a}{\lambda}\right)+\frac{\zeta^{2}}{8}\,\Theta_{2}\left(\frac{a}{\lambda}\right)\right]$ (10)
The modification functions
$\Theta_{1}(\xi)$ $=1-\frac{1}{3}\left(1+0.072\,\xi^{1.13}\right)^{-1},$ (11)
$\Theta_{2}(\xi)$ $=1-\left(1+0.085\,\xi+0.02\,\xi^{0.1}\right)^{-1}$ (12)
depend on the particle radius ($a$) and the DL thickness ($\lambda$). Expanding $\lambda$ using the Gouy–Chapman theory ($\lambda=\sqrt{\frac{k_{B}T\varepsilon}{2N_{A}\mathcal{C}e^{2}}}$), where $\mathcal{C}$ is the molar concentration (i.e., $c^{*}\times 10^{3}$ mol/m3), reveals that $\lambda$, and hence the modification functions, are functions of the local solute concentration. As concentration increases, the Debye layer becomes thinner and phoretic mobility correspondingly increases. The figures below present simulations in S-O and S-I modes, wherein solute concentrations span a broad range ($0.1-10$ mM). Specifically, in the S-O mode, the ambient concentration is set to 0.1 mM while the initial pore concentration is 10 mM; the reverse configuration is used for the S-I mode. For a 250 nm colloid, Fig. A1 (a) shows the variation of mobility across the range of concentrations considered. For a smaller particle, the gap between the (ideal) constant mobility and variable mobility is larger (Fig. A1(d)). In Figs. A1 (b,c), we find that the constant mobility model overestimates the colloidal withdrawal and injection. Figs. A1 (e,f) show how the DL effects further dominate for smaller sized colloids. Fig. A2 compares the two modes for thick DL model. Fig. A2 (a) shows the spatiotemporal variation of mobility in the pore; we note that the mobility rises with pore depth for S-O mode as the concentration increases with $x$, whereas a decrease is noted for S-I mode. Comparing the y-axes, we also note that, beyond $t^{*}\gtrsim t_{d}$, the mobility for the S-I mode does not decay significantly in space, as mobility varies weakly with increasing concentration. This can also be interpreted from Fig. A1(a), where mobility variations in the vicinity of $c^{*}=0.1$ mM are more pronounced, whereas they are much weaker in the vicinity of $c^{*}=10$ mM. Figs. A2(b,c) show that the spatiotemporal patterns of withdrawal and injection follow trends qualitatively similar to those observed in the constant mobility model. Figs. A2(d–f) present analogous results, with the thick DL effect further damping the injection and withdrawal patterns.
In the context of performance, Fig. A3(a,d) shows that the persistence time exhibits a trend similar to that of the constant mobility model. It can also be observed that the persistence in the S-O mode is reduced, this is because of substantial decrease in mobility, particularly near the pore mouth, as seen in Fig. A2(a,d). Additionally, the gradual decline in concentration over time further lowers the mobility. In contrast, for the S-I mode, persistence is slightly enhanced, as the rising concentration over time leads to an increase in mobility, consistent with Fig. A2(a,d). Fig. A3 (b,e), (c,f) show that, relative to the constant mobility model, the thick DL effects dampen the effectiveness trends of withdrawal and injection, respectively. It is interesting to note a finer aspect of withdrawal effectiveness for extremely large solute gradients: in the initial stages, the S-I mode can exhibit slightly enhanced effectiveness compared to the S-O mode, likely due to the nature of spatial pattern at early times. For the S-I mode, the spatial non-monotonicity of $\bm{u}_{p}$ renders near-pore-mouth accumulation, which readily convects and diffuses out due to large curvature. For thick DL mobility model, this brief dominance of the source–injection (S–I) mode at early times is attenuated due to reduced mobility.
Associated Content
Supporting Information: Validation of results and further information on constant charge model, diffusio-osmotic flow, and quantification of distribution spread during colloidal injection.